OpenStudy (anonymous):

px + qy = r 2px - qy = 2r When solving this system of equations, x =

OpenStudy (anonymous):

px+qy=r 2px-qy=2r ................... 3px =3r x=3r/3p=r/p x=r/p

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Np Anytime

OpenStudy (anonymous):

Do you know this one?? Input in standard form the equation of the given line. The line that passes through (-2, 4) and is parallel to x - 2y = 6